OpenStudy (anonymous):

The length of one of the legs of a right triangle is 5 and the hypotenuse is 10. What is the length of the other leg?

OpenStudy (johnweldon1993):

 We can use the Pythagorean theorem ... It states that the sum of each leg squared...= the hypotenuses squared \[a^2 + b^2 = c^2\] where a and b are the legs...and c = the hypotenuse.... so...since we know the hypotenuse...and 1 leg...we can rearrange this to solve for either a or b \[a^2 = c^2 - b^2\] or to solve for just a (what we want) \[a = \sqrt{c^2 - b^2}\] plug in 10 for 'c' and 5 for 'b' and solve for a

OpenStudy (anonymous):

is the answer 75 @johnweldon1993

OpenStudy (johnweldon1993):

That the first part of it...that is what you get when you do c² - b² 10² - 5² 100 - 25 = 75 But remember....this answer is under a square root sign...

OpenStudy (johnweldon1993):

so your final answer would be achieved by answering what is the square root of 75...?